Fermi National Laboratory
 
Operation Summary for December 19, 2011 - December 21, 2011
 

Monday December 19

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, a Muon Rings person began taking beam for studies at 8:59 AM. A Linac RF station (LRF1) tripped off; it reset. Operations established beam to Meson FTBF experiment T-1015 at 10:44 AM. A Booster RF station (BRF4) tripped off; it reset. LRF5 tripped off for the first of two times in 22 minutes; it reset.

During the evening shift, a Booster power supply (B:MP03) alarmed on current regulation; Booster personnel investigated. Muon Rings personnel completed their studies at 5:14 PM. Operations disabled the T-1015 beam at 10:02 PM.

Tuesday December 20

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, the MiniBooNE beam permit tripped off at 12:16 AM due to bad beam positions; it reset. A MI kicker (KPS4DS) tripped off for the first of many times in about 12 minutes on a Thyratron delay fire indication; Operators turned off this kicker and all the gap clearing kickers for the night. This halted sending protons to the Recycler. The problem will be investigated in the morning. Operators completed an MI injection phase tune at 6:57 AM. Operations turned off beam to all the accelerators at 7 AM in preparation for a six-hour access. Operators had Booster, MI, and F-sector ready for accesses and safety system tests by 7:38 AM.

During the day shift, Operators reported that the I- Source was in access at 8:15 AM. Operators had NuMI prepared for access by 8:50 AM. Operators had the Transfer Hall prepared for access by 10:01 AM. Linac personnel reported at 10:26 AM that they had completed their work. Controls personnel reported at 10:56 AM that they had replaced the NuMI BPM VME crates (NBPM6N and NBPM65). Operators had MiniBooNE prepared for access by 11:05 AM. Booster personnel came out of access at 11:27 AM. At 12:17 PM, a Linac RF station (LRF2) tripped off for the first of two times in 11 minutes; it reset. Operations established beam through Linac to the Booster dump at 12:32 PM. At 12:58 PM, MI personnel reported that one of the gap clearing kicker magnets had an open load and needed to be replaced, which would take about three or four hours.

During the evening shift, MI personnel came out of access at 5:15 PM. Operations had beam to NuMI and MiniBooNE by 5:50 PM. Operations established beam to T-1015 at 6:18 PM. At 7:04 PM, a MI/NuMI Lambertson magnet power supply (I:LAM60) tripped off due to excessive ripple and wouldn’t reset; Operators called in a Power Supply technician. Operations loaded protons into the Recycler at 7:17 PM. The Power Supply tech working on I:LAM60 remembered that this had happened before and asked operators to restore crate #73; this restored the ramp and fixed the ripple. At 10:18 PM, BRF11 tripped off for the first of three times in about an hour; it reset. At 11:30 PM, Operators had to reboot a NuMI BPM (NBPM6N).

Wednesday December 21

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operations disabled the T-1015 beam at 12:02 AM. BRF11 tripped off for the first of six times on a ground fault indication at 12:04 AM; it reset. At 12:11 AM, Operators had to reboot NBPM6N. Operators turned off all the devices in the Meson M01 – M05 areas supplied by the MS2 water skid at 4:08 AM; Water Group personnel will add glycol to this system later in the morning. A Muon Rings person began a study at 6:33 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo approved of the plans by Booster and Muon Rings personnel to conduct studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/21/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory